OpenStudy (anonymous):

Will medal and fan!!Gentry poured 23.5 ounces of milk and 7.75 ounces of cream in a mixing bowl. How much more cream could Gentry add to the bowl if the capacity of the bowl is 35.5 ounces? oz

OpenStudy (anonymous):

@Liam2004 @Study312 @heart_offire

OpenStudy (lolacole12):

Notice that because the capacity of the bowl is 35.5 ounces, but there has already been 31.25 ounces total added to it (23.5 ounces + 7.75 ounces), all that needs to be done is subtract 31.25 from 35.5: 35.5 - 31.25 = 4.25 ounces of cream could be added

OpenStudy (anonymous):

4.25 since 31.25 + 4.25 = 35.5 simple as that

OpenStudy (lolacole12):

medal and fan

OpenStudy (anonymous):

lolacole12 lets both medal each other

OpenStudy (lolacole12):

both is correct and ok

OpenStudy (anonymous):

how are both correct...

OpenStudy (lolacole12):

4.25

OpenStudy (anonymous):

since we both said the correct answer it is just that we said it in different ways

OpenStudy (lolacole12):

Liam2004 Best Response Medals 1 4.25 his answer and mine35.5 - 31.25 = 4.25 same

OpenStudy (lolacole12):

both 4.25